OpenStudy (anonymous):

Does anyone now the table for the polar equation -6t(t is degree or radians)? No any helpful websites to help me get the table for this

OpenStudy (richyw):

what are you asking?

OpenStudy (anonymous):

I am asking for the table for the polar equation -6t

OpenStudy (tkhunny):

There is no such thing as "the table". This explains why you cannot find it. Can you provide more information? Is there a specific problem you are working on?

OpenStudy (richyw):

you want to make a table of cartesian coordinates for the polar function \(r(t)=-6t\) ?

OpenStudy (anonymous):

@richyw yes

OpenStudy (tkhunny):

Notice richyw's use of the phrase "a table". You can make a table on any values you like. This is why "the table" doesn't exist.

OpenStudy (anonymous):

Okay. So I am just basically looking for the coordinates of this equation

OpenStudy (richyw):

you could make a table of the polar coordinates. so for whatever angle t, you would have whatever radius r. or you could convert these coordinates to cartesian values, which would have x and y values being compared.

OpenStudy (richyw):

either way all you are doing is solving this equation for a few values.

OpenStudy (anonymous):

No I have polar graph paper and all i need is like the points.

OpenStudy (richyw):

ok I see. so then you just directly solve this. choose whatever values of t that you want and then you will get the r value directly from the equation

OpenStudy (anonymous):

Can you give me an example? If you don't mind

OpenStudy (richyw):

alright well the first point I would do would be at an angle of 0. so\[r(0)=-6(0)=0\] so you know that point will have angle 0 radius 0 (the origin)

OpenStudy (richyw):

how many points do you want?

OpenStudy (anonymous):

I am going by 30 degrees so until 360

OpenStudy (richyw):

are you sure this is in degrees? usually I would assume radians unless it specified.

OpenStudy (anonymous):

well radians will be find also

OpenStudy (richyw):

not that is makes a difference for the shape though

OpenStudy (anonymous):

I understand but I am just trying to graph it

OpenStudy (richyw):

ok well then go 30 degrees more (\(\pi/3\) radians) plug it in \(r(\pi/3)=-6(\pi/3)\) so \(r=-2\pi\)

OpenStudy (richyw):

then plug in \(2\pi/3\) etc until you reach \(2\pi\)

OpenStudy (sirm3d):

uhm, 30 degrees is \(\pi /6\)

OpenStudy (richyw):

haha oops yup

OpenStudy (anonymous):

Your fine you did 60 degrees

OpenStudy (richyw):

yup do what I said but every \(\pi/6\)

OpenStudy (richyw):

I make that mistake so often...

OpenStudy (anonymous):

It's fine we can't remember everything

OpenStudy (richyw):

what I would do if I was you is calculate \(2\pi\) first and then it will give you an idea of the scale you need to use for the graph. because it's going to be a spiral shaped thing

OpenStudy (anonymous):

Yes I have a art math project I already drew my flower

OpenStudy (anonymous):

Hey can you convert that to radians because it will make my day much easier

OpenStudy (richyw):

that's already in radians

OpenStudy (anonymous):

it is

OpenStudy (richyw):

yes if it was in degrees you would have much larger values of r!

OpenStudy (anonymous):

yes that's true

OpenStudy (richyw):

I suppose if you wanted to see the radius scale in terms of fractions of pi I could open up octave. but you are gonna need to make the table of values anyways if that's what's required...

OpenStudy (anonymous):

it is

OpenStudy (sirm3d):

here's a graph of the polar equation \(r(t) = 6t\) in the interval \(0 \leq t \leq 2\pi\)

OpenStudy (anonymous):

Thanks